A car is travelling at a constant speed of 525m/s. Passes a trooper hidden behind a Billboard one second of the car passes the Billboard, the trooper sets out from the Billboard to catch it, acelerating at a constant speed of 50m/s².how long does it take her to take the car?
1
Expert's answer
2020-02-18T09:58:15-0500

Let the time of motion of the trooper is t, then distance traveled by her


"d=\frac{at^2}{2}."

The distance traveled by a car


"d=v(t+1)."

Hence


"\frac{at^2}{2}=v(t+1),"

"\frac{50t^2}{2}=525(t+1),"

"t^2-21t-21=0,"

"t=22\:\rm s."

Finally, the distance traveled by trooper


"d=\frac{at^2}{2}=\frac{50\times 22^2}{2}=12100\:\rm m."